# Discovering reduced order model equations of many-body quantum systems using genetic programming: a technical report

### v1

@@ III Results
Figure 5: Same as Panel a) of Figure 4 but for the case of the neutron density $\rho_{n}(x)$ in DFT for 16 values of $\alpha$ in the testing set. The vertical grey solid line represent the typical spherical DFT calculation time in fortran. This time is faster than for the simpler case in Figure 4 since research codes, like the one used here for DFT, have been optimized for performance. Yet, all SML models which show similar times to those obtained in the case of the modified Gross-Pitaevskii equation are faster by almost 4 orders of magnitude. The horizontal dashed black line shows the median of the optimal coefficients. We observed very similar results for the proton density $\rho_{p}(x)$.
@@
Once the models were trained, we tested their generalization ability on the previously unseen combinations of parameters belonging to the respective test sets. Given a set of parameters $\alpha_{i}$ from the test set, we quantify the accuracy performance as the root mean squared error (or $L^{2}$ norm) between the density obtained from a high-fidelity solver (HF) and the one obtained from the reduced coordinates equation $\phi(x;\alpha_{i})\approx\sum_{k}^{n}\hat{a_{k}}(\alpha_{i})\phi_{k}(x)$, where $\hat{a_{k}}$ were predicted by the respective SML models as a function of $\alpha_{i}$. We quantify the speed performance as the time it takes for each procedure to estimate the reduced coordinates given the target $\alpha_{i}$. Figure 4 shows the performance of each SML model (linear regression and GP) for the modified Gross-Pitaevskii equation. Panel a) shows the accuracy and speed of the three methods. The horizontal dashed black line shows the median error obtained when the coefficients in Eq. (8) are optimized to directly reproduce the test solutions, which serves as a lower bound for the error, only possible to be improved by increasing the number of kept principal components beyond $n=3$. Both GP and LR3-L2 attain residuals comparable with this bound. In terms of speed, all three methods are more than 4 orders of magnitude faster than the HF solver.
@@
Panels b) and c) of Figure 4 show how the first coefficient of the reduced basis $a_{1}$ evolves as the parameters $\kappa$ and $\sigma$ change, respectively, while the other parameters are held constant. This is known as ceteris paribus or isolated effect and is frequently used when interpreting the coefficients of a linear regression model, as it assesses the effect of a particular input feature (predictor) on the response (dependent variable) while keeping all other predictors fixed. Analogous to panel a), the black dashed line represents the evolution of the optimal value of $a_{1}$ as the parameters $\alpha$ change. The GP has appreciably better performance in reproducing this optimal coefficient that the other two methods, and it excells particularly in extrapolating beyond the dashed vertical line which denotes the maximum value of the parameters to which all the models were trained. This suggests that the GP has learned the reduced dynamics with less overfitting.
@@
Figure 5 shows analogous results as panel a) of Figure 4 for the DFT neutron density $\rho_{n}(x)$ of 48Ca (comparable results were obtained for the proton density). Similar to the performance shown in the case of the modified Gross-Pitaevskii equation, the GP method (as well as LR3-L2) is able to learn the dynamics almost perfectly, showing accuracy that is almost on top of the optimized coefficients indicated by the black dashed line. The computation times of all SML methods are comparable to the easier problem in Figure 4. Indeed, it is remarkable that GP is able to find effective equations that explain the change in the nuclear densities as the two parameters $J_{\text{sym}}^{\text{NM}}$ and $L_{\text{sym}}^{\text{NM}}$ are widely varied. The reduced order model we have constructed not only is agnostic to the high dimensional grid-space in the coordinate $x$, but it is also learning the solution manifold of the densities without having to directly track the underlying wave functions, and in the case of DFT the other densities involved. These results poise GP as a promising tool for drastically reducing the computation cost of existing complex nuclear calculations, as well as for model discovery directly from experimental data.

### v2

@@ III Results
Figure 5: Same as Figure 4(a), but for the case of the neutron density $\rho_{n}(x)$ in DFT for 16 values of $\alpha$ in the test set. The vertical grey solid line represents the typical spherical DFT calculation time in FORTRAN. This time is faster than for the simpler case in Figure 4 since research codes, like the one used here for DFT, have been optimized for performance. All supervised machine learning models applied show similar times to those obtained in the case of the modified Gross-Pitaevskii equation, faster by almost 4 orders of magnitude. The horizontal dashed black line shows the median of the optimal coefficients. We observed very similar results for the proton density $\rho_{p}(x)$.
@@
Once the models were trained, we tested their generalization ability on the previously unseen combinations of parameters belonging to the respective test sets. Given a set of parameters $\alpha_{i}$ from the test set, we quantify the accuracy performance as the root mean squared error (or $L^{2}$ norm) between the density obtained from a high-fidelity solver (HF) and the one obtained from the reduced coordinate equation $\phi(x;\alpha_{i})\approx\sum_{k}^{n}\hat{a_{k}}(\alpha_{i})\phi_{k}(x)$, where $\hat{a_{k}}$ were predicted by the respective supervised machine learning models as a function of $\alpha_{i}$. We quantify the speed performance as the time it takes for each procedure to estimate the reduced coordinates given the target $\alpha_{i}$. Figure 4 shows the performance of each SML model (linear regression and GP) for the modified Gross-Pitaevskii equation. Figure 4(a) shows the accuracy and speed of the three methods compared. The horizontal dashed black line depicts the median error obtained when the coefficients in Eq. (8) are optimized to directly reproduce the test solutions, which can serve as a lower bound for the error, only reducible by increasing the number of kept principal components beyond $n=3$. Both GP and LR3-L2 attain residuals comparable to this bound. In terms of speed, all three methods are more than 4 orders of magnitude faster than the HF solver (vertical gray line).
@@
Figures 4(b) and (c) show the dependency of the first reduced basis coefficient $a_{1}$ from parameters $\kappa$ and $\sigma$, respectively, when the other parameters are held constant. This is known as ceteris paribus or isolated effect and is frequently used when interpreting the coefficients of a linear regression model, as it assesses the effect of a particular input feature (predictor) on the response (dependent variable) while keeping all other predictors fixed. As in Figure 4(a), the black dashed line represents the dependency of the optimal value of $a_{1}$ on the parameters $\alpha$. As can be seen, GP has an appreciably better performance in reproducing this optimal coefficient compared to the other methods. Particularly, it excels in extrapolating beyond the dashed vertical line which depicts the maximum parameter values used for the training of models. This suggests that GP has learned extrapolate the reduced dynamics with less overfitting.
@@
Figure 5 shows analogous results to Figure 4(a) for the DFT neutron density $\rho_{n}(x)$ of 48Ca (comparable results were obtained for the proton density). Similar to the performance shown in the case of the modified Gross-Pitaevskii equation, the GP method (as well as LR3-L2) is able to learn the dynamics almost perfectly, showing accuracy that is almost on top of the optimized coefficients indicated by the black dashed line. Computation times for all SML methods are comparable to those in the easier problem of Figure 4. Indeed, it is remarkable that GP is able to find effective equations that explain the change in the nuclear densities as the two parameters $J_{\text{sym}}^{\text{NM}}$ and $L_{\text{sym}}^{\text{NM}}$ are widely varied. The reduced-order model we have constructed not only is agnostic to the high dimensional grid-space in the coordinate $x$, but it is also learning the solution manifold of nuclear densities without having to directly track the underlying wave functions, and in the case of DFT the other densities involved. These results poise GP as a promising tool for drastically reducing the computation cost of existing complex nuclear calculations, as well as for direct model discovery from experimental data.